A deck of playing cards, what is the probability of drawing either a black card or a
6 in a single draw?
1
Expert's answer
2020-11-24T04:34:41-0500

We have 52 cards

Number of ways to draw a black card or 6 = number to draw a black card + number to draw 6 – number to draw black 6

There are 26 black cards (13 clubs and 13 spades)

There are 4 six and 2 of them are black

Therefore we have:

Number of ways to draw a black card or 6 = 26 + 4 – 2 = 28

P(drawing a black card or 6) = "\frac{28}{52}=\frac{7}{13}"